1.sketch the gradient field for the differential equation dy/dt=t-y2 for t=-1......,1

y=-1,.......,1

2.The yield y(t) (in bushes)per acre of a corn crop satisfies the equation dy/dt+y=100+e-t

if y(0)=0 find y at any time t



1
Expert's answer
2021-07-08T13:05:50-0400

Part 1



Part 2

"y'+f(x)y=g(x)" as "I(x) = e^{\int f(x)dx}\\"

Given that the yield y(t) of a corn crop satisfies the equation "\frac{dy}{dt}=100+e^{-t}-y"

"\frac{dy}{dt}=100+e^{-t}-y\\
\frac{dy}{dt}+y=100+e^{-t}\\"

Calculating the integrating factor

"I(t) = e^{\int f(t)dt} = e^t"

Multiplying both sides

"e^t\frac{dy}{dt}+ye^t=100e^{-t}+1\\
(e^ty)'=100e^t+1\\
\int(e^ty)'dt=\int(100e^t+1)\\
e^ty=\int100e^tdt+\int1dt\\
e^ty=100e^t+t+C\\
y= 100+te^{-t}+Ce^{-t}\\
0= 100+0*e^{0}+Ce^{0}\\
C=-100\\
y= 100+te^{-t}-100e^{-t}\\"